OpenStudy (anonymous):

Determine the constant of variation for the direct variation given. (0, 0), (3, 12), (9, 36) 3 4 12

OpenStudy (anonymous):

3

OpenStudy (accessdenied):

Direct variation has the form: \(y = kx\) So, \(k = y/x\) for \(x \ne 0\) Try dividing the y by the x in the nonzero points

OpenStudy (accessdenied):

and no, its not 3. 12/3 = 4 36/9 = 4

OpenStudy (anonymous):

SO THEN.. which one is it?